# Task Singular Vectors: Reducing Task Interference in Model Merging

### v2

@@ 1 Introduction
The widespread availability of pre-trained models and public repositories has driven the development of techniques for efficiently combining and reusing existing models. Among these techniques, model merging approaches enable the creation of multi-task models without additional training. One popular approach, Task Arithmetic (TA) [22], stands out for its simplicity and effectiveness. However, by treating entire networks as high-dimensional vectors, TA and subsequent works overlook crucial structural information [48, 50, 24, 37, 31, 9, 10]. This flattened view limits these approaches to coarse-grained measures like cosine similarity for assessing inter-task interactions.
@@ 2 Related Work
offers an efficient alternative to ensembling by combining existing models without the need for further training. Several approaches address this by determining the neuron permutations that align the models into a shared optimization basin, after which the models are combined using a straightforward averaging method. [1, 25, 41, 35, 8]. An alternative line of approaches focuses on the multi-task scenario, where a single pre-trained model is fine-tuned for different tasks [22, 48, 50, 31, 45, 10, 43, 51]. The prerequisites for methods in this category are outlined in Table 1. Task Arithmetic [22] introduces the concept of task vectors, which are the weight differences between fine-tuned models and the pre-trained base model. By averaging these vectors, a merged multi-task model can be created; conversely, negating a task vector allows for forgetting a specific task. TIES [48] addresses redundancies in model parameters by first selecting the top-$k$ most significant parameter changes and then constructing a sign vector based on the majority sign across all models. The latter is used to merge the task vectors disjointly, meaning the average is not computed when a parameter is zero or when parameters disagree in sign. Similarly, DARE [50] randomly resets redundant parameters to their pre-trained values and rescales the remaining parameters by a factor proportional to the dropped ones, aiming to reduce interference among tasks. Fisher Merging [31] and RegMean [24] merge models by performing weighted averaging, utilizing the Fisher information matrix and the inner product of input vectors, respectively. Model Soups [45] improves performance by averaging the weights of multiple fine-tuned models obtained from a hyperparameter sweep, rather than selecting a single best model. Model Breadcrumbs [10] focuses on merging only significant weights by discarding outliers and both minor and large perturbations in the fine-tuned parameters. More recently, Wang et al. [43] observed that tasks often utilize non-overlapping sets of weights. They propose two methods: the first, TALL-Mask, uses binary masks to activate important task-specific weights, requiring extra storage (as does our method TSV-C) for the masks. The second, Consensus Merging, leverages these masks to remove parameters that are important to fewer than two tasks.
@@ 3.1 Background
We build on Task Arithmetic (TA), which defines task vectors capturing the specific differences in model weights for individual tasks. Formally, the weights $\theta_{\text{MT}}$ of a multi-task model for $T$ tasks are computed by aggregating the task-specific weight differences, or task vectors, as follows:
@@
where $\theta_{\text{pre}}$ is the set of pretrained model weights, $\alpha$ is a scaling factor, and $\tau_{i}=\theta_{\text{ft}_{i}}-\theta_{\text{pre}}$ is the task vector for task $i$, with $\theta_{\text{ft}_{i}}$ being the fine-tuned weights for the task. Differently from TA, however, we consider these operations at the layer level. From a layer-wise perspective, Eq. 1 can be rewritten as
@@
where $\theta_{\text{pre}}^{(l)}$ encodes the pretrained weights for layer $l$, and
$\Delta_{i}^{(l)}=\theta_{\text{ft}_{i}}^{(l)}-\theta_{\text{pre}}^{(l)}$ (3)
is the task-specific weight difference for task $i$ at layer $l$. When layer $l$ has a matrix structure, we call its corresponding $\Delta_{i}{{}^{l}}$ the per-layer task matrix for task $i$. When this is not the case, our framework defaults to standard TA. For brevity, we will generally omit the layer index and refer to the layer-$l$ task matrix $\Delta_{i}^{l}$ as $\Delta_{i}$.
@@
Treating layer-wise weights as structured entities rather than flattened vectors enables us to analyze their SVD, revealing low-rank properties and deeper insight into the inter-task interactions. Given two tasks $i,j$, we consider the SVD of the corresponding task matrices $\Delta_{i}$ and $\Delta_{j}$ at a generic layer:
@@ 3.2 Low-rank nature of layer task matrices
Figure 2: Mean absolute accuracy of the ViT-B-32 model across increasing fractions of retained singular components, averaged over 20 tasks. The red line represents the average accuracy of the original fine-tuned models with full-rank task matrices, while the blue line shows the accuracies using low-rank approximations.
@@ 4 Approach
We demonstrate below how TSVs can be used for both compression and task interference reduction by adapting Eq. 6 for each setting.
@@ 4.1 TSV for compression
This formula yields a low-rank approximation of $\Delta_{h}$, where only the top $k$ singular components of the task-specific matrix are retained. Here, the same number of components $k$ is taken from each task to ensure that each layer is compressed by a factor of $\frac{1}{T}$ relative to the original dimension. Increasing $k$ enhances approximation accuracy but reduces the compression factor.
@@ 4.2 TSV for model merging
10:   $U=P_{U}D_{U}Q_{U}^{\top}$
11:   $V=P_{V}D_{V}Q_{V}^{\top}$
12:  Obtain the orthogonal matrices:
13:   $U_{\bot}=P_{U}Q_{U}^{\top}$
14:   $V_{\bot}=P_{V}Q_{V}^{\top}$
15:  Reconstruct the merged matrix:
16:   $\hat{M}\leftarrow U_{\bot}\Sigma V_{\bot}^{\top}$
17:  Construct merged model weights:
18:   $\theta_{\text{MT}}\leftarrow\theta_{\text{pre}}+\alpha\hat{M}$
19:  return $\theta_{\text{MT}}$
@@ 5.2 Compression results
Our results demonstrate that TSV-C effectively balances compression and performance by leveraging the most significant singular vectors. Maintaining near-original accuracy levels, our approach is particularly suitable for scenarios where storage constraints are critical but high model performance is still required.
@@ 6 Analysis
Figure 5: Approximation error originating by the orthogonalization of the task singular vectors through Procrustes for the ViT-B-32 model across 8 tasks. The violin plots represent layer-wise approximation error distributions for $U$ and $V$ in both full-rank and low-rank cases.
@@ 6.1 Ablation study
TSV-M combines low-rank approximation with task interference reduction. To evaluate the contribution of each component, we conducted an ablation study summarized in Table 4. Applying low-rank approximation alone to the layer task matrices results in worse performance than vanilla Task Arithmetic, shown in the first row where neither component is applied. In contrast, applying task interference reduction alone, while keeping full-rank matrices, significantly improves performance, with gains ranging from $+3.1\%$ to $+6.1\%$. However, the best results are achieved only when both steps are combined in our TSV-M method, yielding substantial performance improvements of $+15.8\%$ to $+17.5\%$.
@@
For example, in the ViT-B-32 model with $8$ tasks, the full-rank setting exhibits a wider error distribution with a higher average, indicating greater variability and larger approximation discrepancies. In contrast, the low-rank setting produces a more compact and lower error distribution, suggesting better consistency in approximation across layers. In the following, we prove that this is not specific to the chosen model but a general property of our approach.
@@
The proof can be found in Section C.3.
Intuitively, the theorem asserts that the approximation error introduced by orthogonalization via Procrustes is smaller when applied to the concatenation of truncated matrices, compared to the concatenation of the original full-rank matrices, provided the rank of the truncated matrices satisfies certain conditions.
To better understand the statement of the theorem, consider the case $T=10$, as in our scenario. The term $\frac{T-2\sqrt{T}}{T}=\frac{10-2\sqrt{10}}{10}\geq\frac{1}{3}$. In our case, we choose $k\approx\frac{n}{T}$. For $T=10$, this gives $k\approx\frac{n}{10}$, which is indeed smaller than $\frac{n}{3}$. Therefore, the condition $k\leq\frac{1}{3}n$ is satisfied.
@@ 6.2 Per-layer task interference
We analyze task interference on a per-layer basis using Eq. 7. As shown in Fig. 6, interference is highest in the initial transformer layers and decreases significantly in the deeper layers. This aligns with the understanding that early, generalized layers capture common features across tasks, increasing the potential for conflict, while deeper layers are more specialized for specific tasks, reducing shared representation and interference. For brevity, we grouped layers within each transformer block in our analysis; each “Layer $n$” in Fig. 6 includes two attention matrices and two MLP matrices. A detailed layer-by-layer interference analysis is provided in Fig. 14.
@@ 6.3 Choice of the interpolation coefficient
The aggregation step outlined in Eq. 2 involves a scaling coefficient $\alpha$, typically tuned using a validation set to optimize performance. The value of this coefficient can have a significant impact on the accuracy of the final model. However, as shown in Fig. 7, our approach consistently achieves the best results with $\alpha=1.0$, indicating that no additional scaling is necessary. Although purely empirical, this finding spares further tuning and validation data, thereby enhancing the practicality and efficiency of our method.
Figure 7: Heatmap of best average normalized accuracy for different alpha values. The vertical labels indicate subsets of 8 tasks selected from the 20 available.
@@ 7 Conclusions
Building on the former insight, we introduced TSV-Compress (TSV-C), a model compression algorithm that reduces task vectors to 10% of their original size while retaining 99% of the original accuracy. Unlike existing methods, our approach maintains a constant storage requirement, independent of the number of tasks. We then developed TSV-Merge (TSV-M), combining low-rank approximation with interference reduction to create a novel model merging technique. Achieving normalized accuracy up to 97%, our method offers a storage-efficient and effective alternative to mixtures of experts, ensembles, and multi-task learning.
Our extensive evaluation, including an ablation study, revealed that while interference reduction contributes most to performance improvements, the optimal results are achieved in combination with compression. We also observed that task interference decreases in deeper layers, and importantly, that our method does not necessitate the tuning of a scaling coefficient.
@@ B.2 Hyperparameter Settings
Figure 11: Mean absolute accuracy of the ViT-B-16 model across increasing fractions of retained singular components, averaged over 20 tasks. The red line represents the average accuracy of the original fine-tuned models with full-rank task matrices, while the blue line shows the accuracies using low-rank approximations.
@@ B.3 Storage cost calculation
Figure 12: Mean absolute accuracy of the ViT-L-14 model across increasing fractions of retained singular components, averaged over 20 tasks. The red line represents the average accuracy of the original fine-tuned models with full-rank task matrices, while the blue line shows the accuracies using low-rank approximations.
@@ C.2 Observations
Notice that $\sum_{i=1}^{n}\hat{\sigma}^{2}_{i}=\text{tr}(\widehat{U}\widehat{U}^{\top})=\text{tr}(\widehat{U}^{\top}\widehat{U})$ and $U^{\top}U$ is a $Tk\times Tk$ matrices with diagonal elements equal to one, so $\text{tr}(\widehat{U}^{\top}\widehat{U})=kT$.
@@
$\sum_{i=1}^{n}\sigma_{i}$ $=\sum_{i=1}^{n}\sqrt{\lambda_{i}(\widehat{U}\widehat{U}^{\top})}$ (12)
@@
$k\leq n\frac{T-2\sqrt{T}}{T}$. This concludes the proof. Since $k$ is a positive number the inequaility of meaningful for $T>4$.
@@ D.2.5 Detailed per-layer task interference
We show in Fig. 14 the per-layer task interference, extending the block-level analysis in the main manuscript.
Figure 16: Absolute accuracy of the ViT-B-32 model across increasing fractions of retained singular components, for each task. The red line represents the accuracy of the original fine-tuned models with full-rank task matrices, while the blue line shows the accuracies using low-rank approximations.

### v3

@@ 1 Introduction
The widespread availability of pre-trained models and public repositories has driven the development of techniques for efficiently combining and reusing existing models. Among these techniques, model merging approaches enable the creation of multi-task models without additional training. $\star$ denotes equal contribution. One popular approach, Task Arithmetic (TA) [22], stands out for its simplicity and effectiveness. However, by treating entire networks as high-dimensional vectors, TA and subsequent works overlook crucial structural information [51, 53, 24, 38, 32, 9, 10]. This flattened view limits these approaches to coarse-grained measures like cosine similarity for assessing inter-task interactions.
@@ 2 Related Work
offers an efficient alternative to ensembling by combining existing models without the need for further training. Several approaches address this by determining the neuron permutations that align the models into a shared optimization basin, after which the models are combined using a straightforward averaging method. [1, 25, 43, 36, 8]. An alternative line of approaches focuses on the multi-task scenario, where a single pre-trained model is fine-tuned for different tasks [22, 51, 53, 32, 48, 10, 46, 54, 40]. The prerequisites for methods in this category are outlined in Table 1. Task Arithmetic [22] introduces the concept of task vectors, which are the weight differences between fine-tuned models and the pre-trained base model. By averaging these vectors, a merged multi-task model can be created; conversely, negating a task vector allows for forgetting a specific task. TIES [51] addresses redundancies in model parameters by first selecting the top-$k$ most significant parameter changes and then constructing a sign vector based on the majority sign across all models. The latter is used to merge the task vectors disjointly, meaning the average is not computed when a parameter is zero or when parameters disagree in sign. Similarly, DARE [53] randomly resets redundant parameters to their pre-trained values and rescales the remaining parameters by a factor proportional to the dropped ones, aiming to reduce interference among tasks. Fisher Merging [32] and RegMean [24] merge models by performing weighted averaging, utilizing the Fisher information matrix and the inner product of input vectors, respectively. Model Breadcrumbs [10] focuses on merging only significant weights by discarding outliers and both minor and large perturbations in the fine-tuned parameters. More recently, Wang et al. [46] observed that tasks often utilize non-overlapping sets of weights. They propose two methods: the first, TALL-Mask, uses binary masks to activate important task-specific weights, requiring extra storage (as does our method TSV-C) for the masks. The second, Consensus Merging, leverages these masks to remove parameters that are important to less than two tasks. Like our approach, TwinMerging [31] and SMILE [44] apply SVD to layer task matrices. However, neglecting singular vector interference, they require a router to selectively activate a subset during inference.
@@ 3.1 Background
We build on Task Arithmetic (TA), which defines task vectors capturing the differences in model weights for individual tasks. Formally, the weights $\theta_{\text{MT}}$ of a multi-task model for $T$ tasks are computed by aggregating the task-specific weight differences, or task vectors, as follows:
@@
where $\theta_{\text{pre}}$ is the set of pretrained model weights, $\alpha$ is a scaling factor, and $\tau_{i}=\theta_{\text{ft}_{i}}-\theta_{\text{pre}}$ is the task vector for task $i$, with $\theta_{\text{ft}_{i}}$ being the fine-tuned weights for the task. Differently from TA, however, we consider these operations at the layer level. From this perspective, Eq. 1 becomes
@@
where $\theta_{\text{pre}}^{(l)}$ encodes the pretrained weights for layer $l$, and $\Delta_{i}^{(l)}=\theta_{\text{ft}_{i}}^{(l)}-\theta_{\text{pre}}^{(l)}$ is the task-specific weight difference for task $i$ at layer $l$. When layer $l$ has a matrix structure, we call its corresponding $\Delta_{i}{{}^{l}}$ the per-layer task matrix for task $i$. When this is not the case, our framework defaults to standard TA. For brevity, we will generally omit the layer index and refer to the layer-$l$ task matrix $\Delta_{i}^{l}$ as $\Delta_{i}$.
@@
Treating layer-wise weights as structured entities rather than flattened vectors enables us to analyze their SVD, revealing low-rank properties and deeper insight into the inter-task interactions. Given two tasks $i,j$, we consider the SVD of their task matrices $\Delta_{i}$ and $\Delta_{j}$ at a generic layer:
@@ 3.2 Low-rank nature of layer task matrices
Figure 2: Mean absolute accuracy of the ViT-B-32 model across increasing fractions of retained singular components, averaged over 20 tasks. The red line represents the average accuracy of the original fine-tuned models with full-rank task matrices, while the green line shows the accuracies using low-rank approximations.
@@ 4 Approach
We demonstrate below how TSVs can be used for both compression and task interference reduction.
@@ 4.1 TSV for compression
This formula yields a low-rank approximation of $\Delta_{h}$, where only the top $k$ singular components of the task-specific matrix are retained. Here, the same number of components $k$ is taken from each task to ensure that each layer is compressed by a factor of $\frac{1}{T}$ relative to the original dimension. Increasing $k$ improves approximation but reduces the compression.
@@ 4.2 TSV for model merging
10:   $U=P_{U}D_{U}Q_{U}^{\top}$   $V=P_{V}D_{V}Q_{V}^{\top}$
11:  Obtain the orthogonal matrices:
12:   $U_{\bot}=P_{U}Q_{U}^{\top}$   $V_{\bot}=P_{V}Q_{V}^{\top}$
13:  Reconstruct the merged matrix:
14:   $\hat{M}\leftarrow U_{\bot}\Sigma V_{\bot}^{\top}$
15:  Construct merged model weights:
16:   $\theta_{\text{MT}}\leftarrow\theta_{\text{pre}}+\alpha\hat{M}$
17:  return $\theta_{\text{MT}}$
@@ 5.2 Compression results
Our results show that TSV-C effectively balances compression and performance by leveraging the most significant TSVs. Maintaining near-original accuracy, our approach is particularly suitable for scenarios where storage constraints are critical but high model performance is still required.
@@ 6 Analysis
Figure 5: Approximation error from the orthogonalization of the TSVs through Procrustes for the ViT-B-32 model across 8 tasks. The violin plots represent layer-wise approximation error distributions for $U$ and $V$ in both full-rank and low-rank cases.
@@ 6.1 Ablation study
TSV-M combines low-rank approximation with task interference reduction. To evaluate the contribution of each component, we conducted an ablation study summarized in Table 4. Applying low-rank approximation alone to the layer task matrices results in worse performance than vanilla Task Arithmetic, shown in the first row where neither component is applied. In contrast, applying task interference reduction alone, while keeping full-rank matrices, significantly improves performance, with gains ranging from $+3.1\%$ to $+6.1\%$. However, the best results are achieved only when both steps are combined, yielding substantial performance improvements of $+15.8\%$ to $+17.5\%$.
@@
For example, in the ViT-B-32 model with $8$ tasks, the full-rank setting exhibits a wider error distribution with a higher average, indicating greater variability and larger approximation discrepancies. In contrast, the low-rank setting produces a more compact and lower error distribution, suggesting better consistency in approximation across layers. In the following, we prove that this is not specific to the chosen model but a general property of our approach. The proof can be found in Section C.3.
@@
Intuitively, the theorem asserts that the approximation error introduced by orthogonalization via Procrustes is smaller when applied to the concatenation of truncated matrices, compared to the concatenation of the original full-rank matrices, provided the rank of the truncated matrices satisfies certain conditions. To better understand the statement of the theorem, consider the case $T=10$, as in our scenario. The term $\frac{T-2\sqrt{T}}{T}=\frac{10-2\sqrt{10}}{10}\geq\frac{1}{3}$. In our case, we choose $k\approx\frac{n}{T}$. For $T=10$, this gives $k\approx\frac{n}{10}$, which is indeed smaller than $\frac{n}{3}$, satisfying the condition $k\leq\frac{1}{3}n$.
@@ 6.2 Per-layer task interference
We analyze task interference on a per-layer basis using Eq. 6. As shown in Fig. 6, interference is highest in the initial transformer layers and decreases significantly in the deeper layers. This aligns with the understanding that early, generalized layers capture common features across tasks, increasing the potential for conflict, while deeper layers are more specialized for specific tasks, reducing shared representation and interference. For brevity, we grouped layers within each transformer block in our analysis; each “Layer $n$” in Fig. 6 includes two attention matrices and two MLP matrices. A layer-by-layer analysis is provided in Fig. 14.
@@ 6.3 Choice of the interpolation coefficient
The aggregation in Eq. 2 involves a scaling coefficient $\alpha$, typically tuned using a validation set, that can have a significant impact on the accuracy of the final model. However, as shown in Fig. 7, our approach consistently achieves the best results with $\alpha=1.0$, indicating that no additional scaling is necessary. Although purely empirical, this finding spares further tuning and validation data, thereby enhancing the practicality and efficiency of our method.
Figure 7: Best average normalized accuracy for different alpha values. The vertical labels indicate different sets of 8 tasks.
@@ 7 Conclusions
Building on the former insight, we introduced TSV-Compress (TSV-C), a model compression algorithm that reduces task vectors to 10% of their original size while retaining 99% of the original accuracy. Unlike existing methods, our approach maintains a constant storage requirement, independent of the number of tasks. We then developed TSV-Merge (TSV-M), combining low-rank approximation with interference reduction to create a novel model merging technique. Achieving normalized accuracy up to 97%, our method offers a storage-efficient alternative to MoEs, ensembles, and multi-task learning.
Our extensive evaluation revealed that while interference reduction contributes most to performance improvements, the optimal results are achieved in combination with compression. We also observed that task interference decreases in deeper layers, and importantly, that our method does not require tuning a scaling coefficient.
@@
Acknowledgments
This work was partially supported by PRIN 2022 project 20227YET9B “AdVVent” CUP code B53D23012830006 and by projects FAIR (PE0000013) and SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. This work was moreover partially supported by the project NEREO (Neural Reasoning over Open Data) project funded by the Italian Ministry of Education and Research (PRIN) Grant no. 2022AEFHAZ.
@@ B.2 Hyperparameter Settings
Figure 11: Mean absolute accuracy of the ViT-B-16 model across increasing fractions of retained singular components, averaged over 20 tasks. The red line represents the average accuracy of the original fine-tuned models with full-rank task matrices, while the green line shows the accuracies using low-rank approximations.
@@ B.3 Storage cost calculation
Figure 12: Mean absolute accuracy of the ViT-L-14 model across increasing fractions of retained singular components, averaged over 20 tasks. The red line represents the average accuracy of the original fine-tuned models with full-rank task matrices, while the green line shows the accuracies using low-rank approximations.
@@ C.2 Observations
Notice that $\sum_{i=1}^{n}\hat{\sigma}^{2}_{i}=\text{tr}(\widehat{U}\widehat{U}^{\top})=\text{tr}(\widehat{U}^{\top}\widehat{U})$ and $\widehat{U}^{\top}U$ is a $Tk\times Tk$ matrices with diagonal elements equal to one, so $\text{tr}(\widehat{U}^{\top}\widehat{U})=kT$.
@@
$\sum_{i=1}^{n}\hat{\sigma_{i}}$ $=\sum_{i=1}^{n}\sqrt{\lambda_{i}(\widehat{U}\widehat{U}^{\top})}$ (11)
@@
Notice that this upper bound is tight, indeed the sum of the singular values of $\hat{U}$ must lie within: $\sqrt{kT}\leq\sum_{i=1}^{n}\hat{\sigma}_{i}\leq kT$. The minimum $\sqrt{kT}$ is achieved if all matrices $U_{i}$ are equals, on the other end, the maximum $kT$ is achieved if the $kT$ columns are orthonormal.
@@
$k\leq n\frac{T-2\sqrt{T}}{T}$. This concludes the proof. Since $k$ is a positive number the inequaility of meaningful for $T>4$. ∎
@@ D.2.5 Detailed per-layer task interference
We show in Fig. 14 the per-layer task interference, extending the block-level analysis in Figure 6 in the main manuscript.
D.2.6 Compression analysis
Our experimental results (see Table 3 in main manuscript) demonstrate that TSV outperforms TALL-Mask on the large-scale ViT-L-14 model for 14 and 20 tasks benchmarks, signaling a scaling advantage. With a fixed-budget analysis, we show in Figure 16 that unlike TALL-Mask, which has a fixed requirement defined by model size and number of tasks, we allow flexible compression rates by allowing rank selection. This enables more aggressive compression, as highlighted in the green region in the figure.
Figure 16: Accuracy with varying compression budgets for ViT-L-14 across 14 tasks.
D.2.7 Test-time adaptation
We compare our method with AdaMerging [52] for test-time adaptation. On a subset of 7 tasks from the 8 task benchmark, AdaMerging achieves an accuracy of 85.43%, while our TSV-M attains 88.93%, an improvement of approximately 3.5% without requiring any test-time adaptation. Additionally, when integrating an AdaMerging-style test-time adaptation into our framework, the accuracy increases to 89.87%, demonstrating the complementary benefits of combining TSV-M with test-time adaptation techniques.
Appendix E Theoretical motivations and analysis
E.1 Theoretical foundation - Empirical design
The TSV-C method is grounded in the well-established framework of low-rank approximation for compression (e.g., [12]). Instead, TSV-M is motivated by more empirical foundations: it is designed to achieve noise reduction through low-rank approximation and to eliminate interference via orthogonalization. Low-rank truncation serves to filter out insignificant variations, while orthogonalization ensures that task-specific singular vectors remain independent, preserving individual task performance.
E.2 Heuristic interference measure
Given that a formal definition of interference in model merging is not yet established, we adopt an operational definition: interference is any cross-task interaction that hinders the merging process. Our proposed Singular Task Interference measure is empirically validated by the consistent performance improvements observed when its value is minimized. Furthermore, we examine the relationship between overlapping singular vectors and knowledge sharing. Unlike multi-task learning (MTL), which enables coordinated knowledge sharing through joint training, the independent task-wise finetuning in model merging may evolve in destructive overlaps in the activations, resulting in interference rather than beneficial knowledge sharing. By orthogonalizing the singular vectors, our approach effectively mitigates these overlaps, reducing interference and enhancing the performance of the merged model.
Figure 17: Absolute accuracy of the ViT-B-32 model across increasing fractions of retained singular components, for each task. The red line represents the accuracy of the original fine-tuned models with full-rank task matrices, while the green line shows the accuracies using low-rank approximations.